Complete combustion of 8.10 g of a hydrocarbon produced 26.0 g of CO2 and 8.88 g of H2O. What is the empirical formula for the hydrocarbon? Insert subscripts as necessary.
1
Expert's answer
2019-09-27T04:47:42-0400

CxHy + O2= xCO2 + y/2H2O

8.1. 26. 8.88

12 x 26/44 = 7.2 g (C)

2 x 8.88/18= 0.9 g (H)

7.2+0.9= 8.1

7.2/12 : 0.9:1= 0.6:0.9= 1:1.5 = 4:6

C4H6.